OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = −one halfx − 2 y = 2x − 2 y = −one halfx − 7

OpenStudy (bahrom7893):

Use y-y1=M(x-x1) and y=Mx+b

OpenStudy (bahrom7893):

so if the lines are parallel, they have the same slope, slope of y=2x+2 is M=2 since that line is in the form of y=Mx+b where both M and b are 2.

OpenStudy (bahrom7893):

sooooo y-y1=2(x-x1) and it's passing through (5,3),so your line is: y-3=2(x-5) y=2x-10+3 or y=2x-7

OpenStudy (anonymous):

Thank you!

OpenStudy (bahrom7893):

np